A particle of mass 1 kg moving along y=x+4 with velocity 3 route2 m/s find angular momentum about origin

Grade:11

1 Answers

Gurmeet
6 years ago
Angular momentum= m*v*perpendicular distancePer.distance of line from original=2root2So Angular momentum=1*3root2*2root2 =12